# A Survey on SAR ship classification using Deep Learning

### v1

@@ Abstract
Deep learning (DL) has emerged as a powerful tool for Synthetic Aperture Radar (SAR) ship classification. This survey comprehensively analyzes the diverse DL techniques employed in this domain. We identify critical trends and challenges, highlighting the importance of integrating handcrafted features, utilizing public datasets, data augmentation, fine-tuning, explainability techniques, and fostering interdisciplinary collaborations to improve DL model performance. This survey establishes a first-of-its-kind taxonomy for categorizing relevant research based on DL models, handcrafted feature use, SAR attribute utilization, and the impact of fine-tuning. We discuss the methodologies used in SAR ship classification tasks and the impact of different techniques. Finally, the survey explores potential avenues for future research, including addressing data scarcity, exploring novel DL architectures, incorporating interpretability techniques, and establishing standardized performance metrics. By addressing these challenges and leveraging advancements in DL, researchers can contribute to developing more accurate and efficient ship classification systems, ultimately enhancing maritime surveillance and related applications.
Index Terms:
SAR ship classification, Deep learning, Synthetic Aperture Radar
@@ I Introduction
Our planet is covered by approximately 71% water, which facilitates over 80% of global trade. This vast commercial activity necessitates maritime surveillance, which is a complex research field due to its extensive coverage and diverse monitoring requirements. These requirements encompass irregular migration, piracy, fisheries management, and traffic monitoring. Maritime surveillance plays a crucial role in promoting environmental sustainability by providing real-time data and insights to tackle pollution, reduce emissions, conserve marine ecosystems, and ensure the long-term health of our oceans.
@@
A critical feature of SAR technology is polarization, which refers to the orientation of the electromagnetic waves used in the radar signals. SAR systems can operate in various polarization modes, such as single-polarization (like VV or HH), dual-polarization (such as VV and VH, or HH and HV), and fully-polarized modes (VV, VH, HH, HV). Here, ”V” stands for vertical polarization and ”H” for horizontal polarization, with the first letter indicating the transmitted polarization and the second the received polarization. These different modes provide varying levels of detail and information about the target, influencing the type of analysis and the performance of specific deep learning algorithms. Different acquisition modes, such as single-look or multi-look, further influence image characteristics. Understanding these fundamental aspects of SAR imagery is essential for effectively utilizing the information contained within these datasets.
@@
Due to coverage, availability and benefits of SAR data, researchers were able to use it for several applications such as Oil Spill Detection [1, 2, 3], Ship Detection [4, 5, 6], Ship velocity estimation [7, 8, 9, 10], and Ship classification [11, 12, 13, 14]. In SAR images, the area represented by each pixel varies depending on the specific sensing platform. For example, the Sentinel-1 C-band SAR satellite constellation provides images of Earth’s surface with a spatial resolution of approximately 20 meters. This limited resolution can render vessels as only a few pixels, making object identification challenging without additional information. To address this limitation and achieve an accurate target labeling, when possible, SAR-derived objects are integrated with AIS data [15].
@@
The SAR data are matched with AIS data using several techniques [16] such as time-based matching, range-based matching and feature-based matching. The SAR images are cropped so that each image chip represents a single detected ship, then labelled according to the matched AIS data and, finally, different Machine Learning models are trained to detect and classify ships from unseen SAR images. The output is usually a set of bounding boxes in the large SAR image indicating the presence of the detected ships and their associated categories. This approach to classification has shown good results using machine learning models, and also deep learning (DL) models are being explored to the same purpose.
@@
Ships exhibit a remarkable diversity in types. For example, the FUSAR dataset [17] encompasses 15 primary ship categories and 98 subcategories. While this extensive dataset proves valuable, ship classification remains a challenging task due to several factors. The limited resolution of SAR images often obscures the unique characteristics of different ships. Additionally, their single-channel nature, as opposed to the multichannel images commonly available from optical sensors, limits the use of color-based features, further complicating object classification. Moreover, the prevalence of cargo ships, the most widely used type, introduces a significant class imbalance, making cargo ships easier to be classified compared to less represented vessels. Machine learning models often struggle with ship classification due to the absence of a clear understanding of which features are most relevant for accurate classification.
@@
Conversely, DL approaches offer a promising solution for ship classification challenges: they can autonomously extract relevant features from satellite imagery, thus aiding vessel identification. However, the effectiveness of DL models heavily relies on crucial decisions regarding data selection, augmentation techniques, learning algorithms, and model architectures. This survey provides a comprehensive review of various techniques employed to train DL models for ship classification.
@@
While existing surveys focus on SAR ship detection using DL [5, 4], a dedicated survey on SAR ship classification using DL is currently absent. This survey addresses a critical gap within the scientific community by providing a comprehensive overview of the state of SAR ship classification using DL. It aims to bridge this gap by summarizing the existing literature, elucidating key methodologies, challenges, and advancements in the field. Additionally, the survey offers insights into past developments, current trends, and future directions in SAR ship classification, thus serving as a valuable resource for researchers and practitioners in this specialized domain. The following chapters outline the structure of this survey: Chapter 2 details the methodology employed for performing survey; Chapter 3 presents the survey’s findings resulting in a DL taxonomy; Chapter 4 reviews relevant literature; Chapter 5 explores key insights from the literature review; Chapter 6 access the design consideration for DL SAR Ship classification; Chapter 7 discusses the key issues; Chapter 8 outlines the future; and Chapter 9 concludes the paper.
@@ II-A Research Questions
    RQ2: What are the performance characteristics (e.g., accuracy, precision, recall) achieved by different DL techniques in SAR ship classification?
    RQ3: What is the impact of incorporating advanced techniques such as pyramid features, feature fusion, and generative methods on the performance of SAR ship classification using DL?
@@
    RQ6: Are there any specific datasets or pre-processing techniques commonly used in the studies on SAR ship classification using DL?
@@ II-B Search Query
Search strings are crucial tools for conducting effective research. They help researchers filter through vast amounts of information to identify relevant and accurate sources. By crafting a well-defined search string, researchers can increase the precision of their search results, saving time and effort in the research process. To retrieve accurate and relevant results focused on DL techniques for SAR ship classification, the following search string was formulated:
@@
This search string aims to capture studies that specifically address the utilization of DL or neural network approaches in SAR ship classification tasks and includes variations of terms such as deep learning, neural network, ship classification, vessel classification, ship categorization, synthetic aperture radar, and SAR.
@@ II-C Study Selection
To identify a relevant set of papers for the systematic literature review, the study selection process involved the following steps:
Step 1: Initial Search
An initial search was conducted using the search query on a range of established academic databases and search engines, including Web of Science, IEEE Xplore, Scopus, Springer, Science Direct, ResearchGate and Google Scholar. This comprehensive approach ensured a thorough examination of the existing literature on the topic.
Step 2: Screening
Two independent reviewers screened the search results by examining the titles and abstracts of the identified papers. Inclusion and exclusion criteria were applied at this stage to determine the relevance of each paper to our research topic.
Step 3: Full-Text Assessment
The papers that passed the initial screening were retrieved in full-text format. Two independent reviewers assessed the full texts of these papers to determine their eligibility for inclusion in the systematic literature review. Any disagreements were resolved through discussion and consensus.
Step 4: Data Extraction
The included papers underwent data extraction, where relevant information was systematically extracted. This included details such as author(s), publication year, objectives, methodology, evaluation metrics, and key findings.
Step 5: Quality Assessment
The quality of the included papers was assessed using appropriate quality assessment tools or checklists. This step aimed to evaluate the methodological rigor and reliability of the selected papers.
Step 6: Data Synthesis
Finally, the data extracted from the included papers were synthesized and analyzed to derive meaningful insights, identify trends, and answer the research questions outlined in our systematic literature review.
The study selection process was conducted meticulously to ensure the inclusion of high-quality and relevant papers that contribute to the body of knowledge on SAR ship classification using DL. The inclusion criteria that were applied during the study selection process were inspired by [18], which derives them from the PICOC approach (Population, Intervention, Comparison, Outcome, Context) proposed by [19] to ensure a focused and comprehensive selection.
@@ II-D Inclusion Criteria
This survey includes papers focused specifically on ship classification from satellite SAR images using DL techniques, algorithms or models, and appearing in peer-reviewed journals, conference proceedings or other reputable scholarly sources. Additional criteria to be satisfied were
@@
All the papers not complying with these requirements were excluded from any further analysis.
@@ III-A Search and screening Results
The search query was executed across multiple databases; the papers were retrieved from each database through the mentioned query string, Section II-B, with no further selection based on relevance and inclusion criteria. The numbers of retrieved titles are detailed in Table I.
TABLE I: Number of Papers Retrieved from Different Databases
Database Number of Papers
Web of Science 52
IEEE Xplore 50
Scopus 63
Springer 87
Science Direct 30
Research Gate 100
Google Scholar 1450
The initial screening reduced the total number of papers from 1832 to 161 by excluding duplications, irrelevant search results and papers that fell outside the defined inclusion criteria. Figure 1 illustrates the distribution of papers remaining after this assessment, specifying their provenance by database.
WOSIEEEScopusSDGS0020202020404040406060606080808080100100100100333333332323232319191919222284848484DatabaseNumber of PapersDistribution of Papers Across Databases
Figure 1: Distribution of titles surviving the initial screening.
During the full-text assessment stage, the content of the remaining 161 papers was examined to determine their relevance to research questions and inclusion criteria. The aim of this analysis was to identify the most pertinent studies contributing significantly to the development of DL-based SAR ship classification.
By rigorous evaluation of the full texts of these papers, it was ensured that only high-quality, relevant, and valid studies were included in the final selection. This process is essential for maintaining the integrity and reliability of this systematic literature review and for extracting meaningful insights to address the research objectives. The full-text assessment further refined the set of papers, reducing them to a final number of 49 (see Table II) and allowing for the derivation of well-founded conclusions and evidence-based findings.
TABLE II: Comprehensive List of Relevant Research Papers
Year Study
@@
2024 [65], [66], [67], [68]
After the meticulous and thorough evaluation of the full texts described above, we derived the taxonomy illustrated in Figure 2, which shows models, datasets, augmentations and learning techniques used in the literature for SAR ship classification.
Figure 2: DL Taxonomy in SAR Images for Ship classification: The ”*” in CNN denotes convolutional neural networks that were specifically designed by researchers rather than using predefined architectures.
@@ III-B Survey Findings
The analysis of relevant literature revealed several key insights into the utilization of DL techniques for ship classification on SAR images. Here is a list thereof, ordered by research questions (see Section II):
@@ 1. RQ1
    Evolution of Deep Learning Techniques: The adoption of CNN models initially stemmed from their success in other image processing tasks. However, as the demand for enhanced performance intensified, researchers integrated various image processing techniques and handcrafted features to augment the effectiveness of DL models [31, 68, 67].
@@ 2. RQ2
    Performance Characteristics: Accuracy is the most commonly used performance metric in the literature for evaluation. However, its widespread use highlights the need for incorporating additional metrics, particularly when dealing with long-tailed, imbalanced datasets. Relying solely on accuracy can be misleading in such scenarios, as it may not adequately capture the model’s performance across different classes, especially for minority classes.

### v2

@@ Abstract
Deep learning (DL) has become a central approach for ship classification using synthetic aperture radar (SAR) imagery. This survey reviews 74 representative studies selected from 187 publications, categorizing them into a taxonomy with four main dimensions: (i) DL architectures, (ii) datasets, (iii) image augmentation, and (iv) learning techniques. We analyze how approaches such as handcrafted feature integration, data augmentation, fine-tuning, and transfer learning influence classification performance, and we summarize the use of public benchmarks including OpenSARShip and FUSARShip. The survey highlights key challenges: limited data availability, class imbalance, lack of standardized metrics, and limited interpretability of DL models. Future research directions include the development of SAR-specific DL architectures, advanced augmentation and generative approaches, integration of handcrafted and deep features, interpretable DL, and stronger interdisciplinary collaboration. By addressing these challenges, DL-based SAR ship classification can achieve greater robustness, accuracy, and transparency, ultimately strengthening maritime surveillance and operational monitoring.
@@ I Introduction
Our planet is covered by approximately 71% water, which facilitates over 80% of global trade. This vast commercial activity necessitates maritime surveillance, which is a complex research field due to its extensive coverage and diverse monitoring requirements. These requirements include irregular migration, piracy, fisheries management, and traffic monitoring. Maritime surveillance also plays a crucial role in promoting environmental sustainability by providing real-time data and insights to tackle pollution, reduce emissions, conserve marine ecosystems, and ensure the long-term health of our oceans.
@@
A critical feature of SAR technology is polarization. SAR systems operate in various polarization modes, such as single-polarization (VV or HH), dual-polarization (VV and VH, or HH and HV), and fully polarized modes (VV, VH, HH, HV). Here, “V” stands for vertical polarization and “H” for horizontal polarization, with the first letter indicating the transmitted polarization and the second the received polarization. These different modes provide varying levels of detail and information about the target, influencing the type of analysis and the performance of specific DL algorithms. Different acquisition modes, such as single-look or multi-look, further influence image characteristics. Understanding these fundamental aspects of SAR imagery is essential for effective utilization of the information contained within these datasets. Due to their broad coverage, availability, and specific benefits, SAR data can be used for several applications, such as oil spill detection [1, 2, 3], ship detection [4, 5, 6], ship velocity estimation [7, 8, 9, 10], and ship classification [11, 12, 13, 14]. In SAR images, the area represented by each pixel varies depending on the specifics of the sensing platform. For example, the Sentinel-1 C-band SAR satellite constellation provides images of Earth’s surface with a spatial resolution of approximately 20 meters. This limited resolution can render vessels as only a few pixels, making object identification challenging without additional information. To address this limitation and achieve accurate target labeling, when possible, SAR-derived objects are integrated with AIS data [15].
@@
To properly generate training sets for classification, SAR data can be matched with AIS data using several techniques [16] such as time-based matching, range-based matching and feature-based matching. Usually, SAR images are cropped so that each image chip represents a single detected ship, which is then labeled according to the matched AIS data. Then, different machine learning models are trained to detect and classify ships from unseen SAR images. The output is usually a set of bounding boxes in the large SAR image, indicating the presence of the detected ships and their associated categories. This approach to classification has shown good results using classical machine learning models, but DL models are also being explored for the same purpose.
@@
Ships exhibit a remarkable diversity in types. For example, the FUSAR dataset [17] encompasses 15 primary ship categories and 98 subcategories. whereas this extensive dataset proves valuable, ship classification remains a challenging task due to several factors. The limited resolution of SAR images often obscures the unique characteristics of different ships. Additionally, their single-channel nature, as opposed to the multichannel images commonly available from optical sensors, prevents the use of color-based image processing. Moreover, the prevalence of cargo ships, the most widely used type, introduces a significant class imbalance, making cargo ships easier to be classified compared to less represented vessels. Machine learning models often struggle with ship classification due to the absence of a clear understanding of which features are most relevant for accurate classification.
@@
Conversely, DL approaches offer a promising solution for ship classification challenges because they can autonomously extract relevant features from satellite imagery, thus aiding vessel identification. However, the effectiveness of DL models relies heavily on crucial decisions regarding data selection, augmentation techniques, learning algorithms, and model architectures. This survey provides a comprehensive review of various techniques used to train DL models for ship classification.
@@
Whereas existing surveys focus on SAR ship detection using DL [5, 4], a dedicated survey on SAR ship classification using DL is currently absent. This survey addresses a critical gap within the scientific community by providing a comprehensive overview of the state of SAR ship classification using DL. It aims to bridge this gap by summarizing the existing literature, elucidating key methodologies, challenges, and advances in the field. This survey also offers insights into past developments, current trends, and future directions in SAR ship classification, thus serving as a valuable resource for researchers and practitioners in this specialized domain. The following sections outline the structure of this paper. Section 2 details the survey methodology; Section 3 presents our findings, resulting in a DL taxonomy; Section 4 reviews the relevant literature; Section 5 explores key insights from the literature review; Section 6 provides design considerations for DL SAR Ship classification; Section 7 discusses the key issues; Section 8 outlines possible future developments, and Section 9 concludes the paper.
Contributions and Scope
Contributions
This survey gives the following contributions:
  • A systematic review of 187 papers, the most pertinent 74 thereof are analyzed in depth. These references are listed in Table II, and their key features are summarized in Section IV.
  • A clear taxonomy (Figure 2) covering (i) DL architectures, (ii) datasets, (iii) image augmentation, and (iv) learning techniques (including transfer learning and fine-tuning).
  • A standardized performance table, reporting scores, classes, and datasets found in the reviewed works, with emphasis on benchmarks such as OpenSARShip and FUSARShip. The performance indices are consistently reported using F1 score and overall accuracy (OA).
  • A synthesis of empirical trends, including the effectiveness of shallow CNNs under data scarcity, the benefits of data augmentation and fine-tuning, strategies for class imbalance (such as data splitting and tailored loss functions), and the impact of feature/polarization fusion, metric learning, multi-scale/pyramid learning, learning rate scheduling, and architecture search (NAS).
  • Practical steps to mitigate data scarcity, including creating high-quality datasets from open-source SAR imagery, establishing AIS–SAR mapping methodologies for reliable labels, and encouraging standardization and data sharing across the community.
  • A selection of open issues, such as data scarcity, class imbalance, standardized reporting, interpretability, and reproducibility.
  • Concrete directions to guide future work in SAR ship classification.
Scope
The focus of this survey is SAR-based ship classification (category-level recognition on cropped ship chips/SAR patches) using DL. We included studies that (i) operate on SAR imagery from spaceborne or airborne platforms, (ii) consider single-, dual-, or fully-polarized data, and (iii) report F1 score and/or overall accuracy (OA), which serve as the basis for comparison in this survey [^1] When available, F1 is treated as macro-F1; if a source reports a different F1 variant, it is indicated in the performance table. Likewise, any form of accuracy reported in the source (e.g., top-1 accuracy, classification accuracy, recognition rate) is consistently referred to as overall accuracy (OA) in this survey.. We include discussion of handcrafted feature baselines where they inform DL design and performance. We excluded studies limited to detection or segmentation without a classification component, optical-only methods without a SAR pathway, and non-maritime Automatic Target Recognition (ATR) unless directly comparable to ship classification.
@@ II-A Research Questions
    RQ2: What are the performance characteristics (e.g., accuracy, F1) achieved by different DL techniques in SAR ship classification?
    RQ3: What is the impact of incorporating advanced techniques (such as pyramid features, feature fusion, and generative methods) on the performance of SAR ship classification using DL?
@@
    RQ6: Are there any specific datasets or preprocessing techniques commonly used in the studies on SAR ship classification using DL?
@@ II-B Search Query
Search strings are crucial tools to filter through vast amounts of information and identify relevant and accurate sources. The search string formulated to capture studies that are specifically part of our scope was
@@
Apparently, this string addresses the use of DL approaches in SAR ship classification, including variations of terms such as DL, neural network, ship classification, vessel classification, ship categorization, synthetic aperture radar, and SAR.
@@ II-C Study Selection
Records identified: 1787 Duplicates removed: 912 Titles/abstracts screened: 875 Excluded: 585 Full-text assessed: 187 Excluded with reasons: 113 Studies included in synthesis: 74
Figure 1: Screening flow.
To identify a relevant set of papers for the systematic literature review, the study selection process (Figure 1) involved the following steps:
1. Initial Search
An initial search was conducted using the search query reported in II-B on a range of established academic databases and search engines, including Web of Science, IEEE Xplore, Scopus, Springer, Science Direct, ResearchGate and Google Scholar. This comprehensive approach ensured a thorough examination of the existing literature on the topic.
2. Screening
Two independent reviewers screened the search results by examining the titles and abstracts of the identified papers. Inclusion and exclusion criteria were applied at this stage to determine the relevance of each paper to our research.
3. Full-Text Assessment
The papers that passed the initial screening were retrieved in full-text format. Two independent reviewers then assessed the full texts of these papers to determine their eligibility for inclusion in our review. Any disagreements were resolved through discussion and consensus.
4. Data Extraction
Relevant information was systematically extracted from the included papers, including details such as author(s), publication year, objectives, methodology, evaluation metrics, and key findings.
5. Quality Assessment
The quality of the included papers was assessed using appropriate tools or checklists. This step aimed to evaluate the methodological rigor and reliability of the selected papers.
6. Data Synthesis
Finally, the data extracted from the included papers were synthesized and analyzed to derive meaningful insights, identify trends, and answer the research questions outlined above.
The study selection process was conducted meticulously to ensure the inclusion of high-quality and relevant papers that contribute to the body of knowledge on SAR ship classification using DL. The inclusion criteria applied during the selection process were inspired by [18], which derives them from the PICOC approach (Population, Intervention, Comparison, Outcome, Context) proposed by [19] to ensure a focused and comprehensive selection.
@@ II-D Inclusion Criteria
This survey includes papers focused specifically on ship classification from satellite SAR images using DL techniques, algorithms or models, and appearing in peer-reviewed journals, conference proceedings or other reputable scholarly sources. Additional criteria to be satisfied were:
@@
All the papers not complying with these requirements were excluded from any further analysis. [^2] Threats to Validity: Selection bias: despite multi-database search, some works may be missed. Metric comparability: heterogeneous splits limit strict cross-paper comparisons. Reporting variance: inconsistent metric definitions (F1 variants) and missing details can affect synthesis. We mitigate these by transparent inclusion criteria, deduplication, and explicit caveats in tables. Reproducibility Checklist: We provide: (i) the performance table (tables-V, VI); (ii) search strings (section-II-B); (iii) deduplication logic summary (table-I). Each entry records dataset, split notes, and metrics (F1/OA).
@@ III-A Search and Screening Results
The papers were retrieved by executing the search query (see Section II-B) across multiple databases, followed by selection based on the predefined relevance and inclusion criteria. Per-database counts after deduplication are reported in Table I. The initial screening reduced the corpus from 1,847 records to 187 by removing duplicates, off-topic items, and studies that did not meet the inclusion criteria.
TABLE I: Search results before/after deduplication and screening.
Source Initial After dedup Included
Web of Science 52 33 13
IEEE Xplore 55 43 18
Scopus 63 19 7
ScienceDirect 87 2 1
SpringerLink 30 - -
Google Scholar 1460 90 35
ResearchGate [^3] Used primarily for discovery; entries were verified against peer-reviewed sources. 100 - -
Total (unique) 1787 187 74
During the full-text assessment stage, the content of the remaining 187 papers was examined to determine their relevance to research questions and inclusion criteria. The aim of this analysis was to identify the most pertinent studies contributing significantly to the development of DL-based SAR ship classification.
By rigorous evaluation of the full texts of these papers, it was ensured that only high-quality, relevant, and valid studies were included in the final selection. This process was essential for maintaining the integrity and reliability of our literature review and for extracting meaningful insights to address the research objectives. The full-text assessment further refined the set of papers, reducing them to a final number of 74 (see Table II) and allowing for the derivation of well-founded conclusions and evidence-based findings.
TABLE II: Chronological Summary of SAR Ship Classification Papers
Year Published Studies
@@
2024 [68], [69], [70], [71], [72], [73], [74], [75], [76], [77], [78], [79], [80], [81], [82], [83]
2025 [84], [85], [86], [87], [88], [89], [90], [91], [92], [93]
After the meticulous and thorough evaluation of the full texts described above, we derived the taxonomy illustrated in Figure 2, which shows models, datasets, augmentations and learning techniques used in the literature for DL-based SAR ship classification.
DL based SAR-Ship Classification Architectures
CNN lineage
— Classic (AlexNet, VGG)
— Residual (ResNet, DenseNet) — Inception — Custom CNN Transformer / Attention
— ViT, SE, CBAM
Temporal / Hybrid
— CNN–LSTM
Generative
— GAN
Datasets
Benchmark collections
— OpenSARShip (OS)
— FUSAR-Ship (FS) Satellite scenes
— COSMO-SkyMed
— TerraSAR-X — PALSAR-2 — GF-3 Mixed / Private
— FS+OS, Author-specific
Image Augmentation
— Geometric (rot./flip/scale …)
— Noise/Spectral (speckle, polar jitter) — Synthetic (CutMix, GAN-gen.)
Learning Techniques
— Feature / Polarisation fusion
— Multiscale processing — Metric & Siamese learning — Attention mechanisms — Domain adaptation / TL — Semi- / Self-supervised — Knowledge distillation — Sampling / Imbalance — Curriculum learning — Super-resolution — Neural arch. search
Figure 2: Taxonomy of DL methods for SAR ship classification. The taxonomy organizes prior work into four dimensions: (i) architectures, (ii) datasets, (iii) image augmentation, and (iv) learning techniques. denotes CNN variants specifically designed by the respective authors.
@@ III-B Survey Findings
The analysis of relevant literature revealed several key insights into the use of DL techniques for ship classification from SAR images. Here is a list thereof, ordered by research questions (see Section II):
@@ 1. RQ1
    Evolution of DL Techniques: The adoption of CNN models initially stemmed from their success in other image processing tasks. However, as the demand for enhanced performance intensified, researchers integrated various image processing techniques and handcrafted features to augment the effectiveness of DL models [31, 71, 70, 81, 63, 94].
@@ 2. RQ2
    Performance Characteristics: Accuracy is the most commonly reported performance metric in the literature. However, additional metrics should be considered, particularly when addressing long-tailed or imbalanced datasets [83]. Relying solely on accuracy can be misleading, as it may not adequately reflect performance across different classes, especially minority categories. To address this, we extracted the reported accuracy and F1 scores from the studies and summarized them in Tables V and VI. It is important to note that these scores are not directly comparable across studies due to differences in dataset selection: subsets of test data vary, and even when the same number of classes is reported, the specific class composition may differ.